If we assume that a spaceship could accelerate from rest at a constant rate of 9.81 m/s2, then how long would it take to reacPLEASE SHOW WORK
Answer #1

speed to achieve = 0.02*3*10^8 = 6*10^6 m/s

acceleration = 9.81 m/s^2

using the equation of linear motion we have

v = u + at

6*10^6 = 0 + 9.81*t

t = 6.11*10^5 s

so fourth option is the correct answer